# LoRA-Edit: Controllable First-Frame-Guided Video Editing via Mask-Aware LoRA Fine-Tuning

### v3

@@ 4.1 Implementation Details
We conduct our experiments using videos consisting of 49 frames, with a resolution of either or . All main results are obtained using the Wan2.1-I2V 480P model. Additional results based on HunyuanVideo-I2V are included in Sec. 4.4. Our framework is built upon the publicly available diffusion-pipe codebase111https://github.com/tdrussell/diffusion-pipe. For each video editing sample, we begin by training on the input video for 100 steps as described in Sec. 3.3.1. If additional edits are applied to later frames, we continue training for another 100 steps on data that includes additional modifications as described in Sec. 3.3.2. This helps the model incorporate user-specified appearance changes. We use a learning rate of for all experiments. Training on 49-frame videos typically requires 20 GB of GPU memory. In Sec. 4.5, we describe a memory optimization strategy that reduces GPU requirements.
@@ 4.2 Comparison with State-of-the-Arts
Comparison with Reference-Guided Video Editing.
We compare our method with two recent reference-guided video editing approaches: Kling1.6 (KlingAI, 2025) and VACE (Jiang et al., 2025). To evaluate the performance on reference-guided video editing, we collect 20 high-quality video clips from Pexels and YouTube. Each video is paired with a reference image representing the desired edit. We use ACE++ (Mao et al., 2025) to apply the edit to the first frame for our method. Figure 4 shows visual comparison results. Compared to Kling1.6 and VACE, our method better respects the intended appearance in the edited region while preserving background content and temporal consistency.
Comparison with First-Frame-Guided Video Editing.
We further compare our method with recent first-frame-guided video editing approaches, including I2VEdit (Ouyang et al., 2024), Go-with-the-Flow (Burgert et al., 2025), and AnyV2V (Ku et al., 2024). All baselines take the edited first frame as input and attempt to propagate the edits through the entire sequence. To ensure a fair and consistent evaluation, we adopt the test set from I2VEdit, which contains videos from diverse sources along with paired first-frame edits. Figure 5 shows qualitative results. In the portrait example (left), our method accurately adds the necklace while preserving the facial structure, while baseline methods often distort the face or produce artifacts. In the street scene (right), our approach transfers the clothing style cleanly across frames without affecting the background, whereas baseline methods distort the clothing or introduce changes in unedited areas.
Quantitative Results.
For quantitative evaluation on first-frame-guided video editing, we use three metrics: 1) DeQA Score You et al. (2025), a state-of-the-art method for assessing image quality; 2) CLIP Score, which measures the semantic alignment between generated frames and edited first frame by comparing their CLIP Radford et al. (2021) embedding similarity; and 3) Input Similarity, which computes the CLIP embedding similarity between the generated frames and the input frames on a per-frame basis. As shown in Tab. 4.2, our method outperforms others across all metrics. For quantitative evaluation on reference-guided video editing, we conducted a user study with 35 participants. Each participant was randomly shown 10 groups of results generated by different methods. For each group, the participants were asked to rank the results based on motion consistency and background preservation. Tab. 4.2 demonstrates the superiority of our method in both aspects.
@@ 4.3 Ablation Studies
Disentangling Edits and Background.
To validate the effectiveness of spatial conditioning in separating edited regions from preserved content, we conduct an ablation study comparing our method with and without spatial condition inputs. Figure 6 shows the results. On the left, the goal is to apply a hair color change. Without spatial conditioning, the edit is applied globally, altering the lighting across the frame. In contrast, with spatial conditioning, the model localizes the change to the hair region while leaving the background untouched. A similar effect is observed in the right example, where clothing edits are confined to the target area only when spatial condition is used.
Appearance Control in Propagated Edits.
We conduct an ablation to evaluate the impact of using edited frames beyond the first frame for controlling appearance in edits propagation. Figure 7 compares two settings: using only the first frame as input versus adding an edited frame at a later timestep. While using only the first frame can still generate reasonable results, incorporating an additional edited frame offers stronger control over the appearance, leading to more consistent and accurate propagation of the intended edit.
4.4 Results based on HunyuanVideo-I2V Model
In addition to the main results based on the Wan2.1-I2V 480P (Wang et al., 2025) model, we also conducted experiments using the HunyuanVideo-I2V (Kong et al., 2024), with a resolution of . Figure 8 presents the results obtained with the HunyuanVideo-I2V model, which performs well in generating temporally consistent edits while preserving the background content. Although the HunyuanVideo-I2V model’s performance is not as strong as Wan2.1-I2V, it still demonstrates the effectiveness of our method across different models.
4.5 Low-Cost Training Strategy
Training on a 49-frame video can be challenging due to the high GPU memory requirements. To address this, we propose a memory-efficient strategy by splitting the video into smaller overlapping clips. Specifically, the 49 frames are divided into four clips: (1) frames 1 to 13, (2) frames 13 to 25, (3) frames 25 to 37, and (4) frames 37 to 49. Each clip contains 13 frames, significantly reducing the VRAM load during training. The results of this approach are shown in Figure 9. While this method helps reduce memory usage, it introduces some artifacts such as vertical stripes. However, for scenes with minimal motion, like the example on the right, these artifacts are barely noticeable. This approach offers a practical solution for training on consumer-grade hardware, with performance similar to training on the entire 49-frame video.
5 Conclusion and Limitation
@@
Despite these advantages, a key limitation is that LoRA fine-tuning requires substantial computational resources and training time. While this is acceptable for most high-precision video editing applications, it may limit accessibility in resource-constrained environments. Future work will focus on developing more efficient LoRA optimization strategies to address this challenge.

### v6

@@ 4.1 Implementation Details
We conduct our experiments using videos consisting of 49 frames, with a resolution of either or . All main results are obtained using the Wan2.1-I2V 480P model. Additional results based on HunyuanVideo-I2V are included in App. E. Our framework is built upon the publicly available diffusion-pipe codebase111https://github.com/tdrussell/diffusion-pipe. Details regarding our automated mask acquisition workflow are provided in App. A. For each video editing sample, we begin by training on the input video for 100 steps as described in Sec. 3.3.1. If additional edits are applied to later frames, we continue training for another 100 steps on data that includes additional modifications as described in Sec. 3.3.2. This helps the model incorporate user-specified appearance changes. We use a learning rate of for all experiments. Training on 49-frame videos typically requires 20 GB of GPU memory. In App. C, we describe a strategy that reduces GPU requirements.
@@ 4.2 Comparison with State-of-the-Arts
Comparison with Reference-Guided Video Editing. We compare our method with two recent reference-guided video editing approaches: Kling1.6 (KlingAI, 2025) and VACE (Jiang et al., 2025). To evaluate the performance on reference-guided video editing, we collect 20 high-quality video clips from Pexels and YouTube. Each video is paired with a reference image representing the desired edit. We use ACE++ (Mao et al., 2025) to apply the edit to the first frame for our method. Figure 4 shows visual comparison results. Compared to Kling1.6 and VACE, our method better respects the intended appearance in the edited region while preserving background content and temporal consistency.
Comparison with First-Frame-Guided Video Editing. We further compare our method with recent first-frame-guided video editing approaches, including I2VEdit (Ouyang et al., 2024), Go-with-the-Flow (Burgert et al., 2025), and AnyV2V (Ku et al., 2024). All baselines take the edited first frame as input and attempt to propagate the edits through the entire sequence. To ensure a fair and consistent evaluation, we adopt the test set from I2VEdit, which contains videos from diverse sources along with paired first-frame edits. Figure 5 shows qualitative results. In the portrait example (left), our method accurately adds the necklace while preserving the facial structure, while baseline methods often distort the face or produce artifacts. In the street scene (right), our approach transfers the clothing style cleanly across frames without affecting the background, whereas baseline methods distort the clothing or introduce changes in unedited areas.
Quantitative Results. For quantitative evaluation on first-frame-guided video editing, we use three metrics: 1) DeQA Score You et al. (2025), a state-of-the-art method for assessing image quality; 2) CLIP Score, which measures the semantic alignment between generated frames and edited first frame by comparing their CLIP Radford et al. (2021) embedding similarity; and 3) Input Similarity, which computes the CLIP embedding similarity between the generated frames and the input frames on a per-frame basis. As shown in Tab. 1, our method outperforms others across all metrics. For quantitative evaluation on reference-guided video editing, we conducted a user study with 35 participants. Each participant was randomly shown 10 groups of results generated by different methods. For each group, the participants were asked to rank the results based on motion consistency and background preservation. Tab. 2 demonstrates the superiority of our method in both aspects.
@@ 4.3 Ablation Studies
Disentangling Edits and Background. To validate the effectiveness of mask conditioning in separating edited regions from preserved content, we conduct an ablation study comparing our method using a foreground-background mask against a baseline version without it. Figure 6 shows the results. On the left, the goal is to apply a hair color change. Without mask conditioning, the edit is applied globally, altering the lighting across the frame. In contrast, with mask conditioning, the model localizes the change to the hair region while leaving the background untouched. Similarly, in the right example, clothing edits are confined to the target area only with mask conditioning. Appearance Control in Propagated Edits. We conduct an ablation to evaluate the impact of using edited frames beyond the first frame for controlling appearance in edits propagation. Figure 7 compares two settings: using only the first frame as input versus adding an edited frame at a later timestep. While using only the first frame can still generate reasonable results, incorporating an additional edited frame offers stronger control over the appearance, leading to more consistent and accurate propagation of the intended edit.
Mask Robustness. To evaluate the sensitivity of our method to mask quality, we conducted an ablation study comparing results under three mask configurations: a tight mask using high-precision segmentation directly from SAM2 (Ravi et al., 2024); a noisy mask simulated by downsampling the segmentation to a grid to introduce significant boundary errors; and the bounding box mask we used, which employs a loose rectangular region derived from the segmentation. The comparisons in Figure 8 reveal a key insight: pixel-perfect precision is often unnecessary and can be restrictive for generative editing. Since the goal is to alter an object’s appearance, the generated entity requires a spatial buffer to undergo necessary contour variations. A strictly tight mask would excessively constrain the generation, forcing the new object to adhere rigidly to the original silhouette and potentially clipping natural details. In contrast, we observe that loose masks including both the highly perturbed “Noisy Mask” and our “Bounding Box” yield better visual results than tight segmentation. By conditioning on a looser region, we allow the model to utilize its strong priors to heal the boundary between the edit and the frozen background. This robustness confirms that our framework relies on the mask for semantic localization rather than strict pixel clipping, validating our design choice of using an automated, approximate masking workflow (See App. A).
5 Conclusion
@@
Acknowledgements
This work is supported in part by the Centre for Perceptual and Interactive Intelligence (CPII) Ltd., a CUHK-led InnoCentre under the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government. This work is also support by RGC Early Career Scheme (ECS) No. 24209224, CUHK-CUHK(SZ)-GDSTC Joint Collaboration Fund No. 2025A0505000053. We would like to thank Yumeng Shi for her valuable suggestions on the figures and demo. We also thank the reviewers for their insightful comments.
Ethics statement
We have read and adhered to the ICLR Code of Ethics. Our work, centered on a controllable video editing framework, aims to advance creative, commercial, and scientific applications. We acknowledge, however, that generative video technologies can be misused for creating misleading or harmful content, such as deepfakes. Our research is intended to provide artists and creators with more flexible and precise tools, not to facilitate malicious use. We advocate for the responsible development and deployment of generative models, accompanied by robust detection mechanisms and clear content provenance standards to mitigate such risks. Furthermore, our method relies on pre-trained foundation models (Wan2.1-I2V, HunyuanVideo-I2V), which may inherit biases from their training data. While our approach does not introduce new data, we recognize that these biases could be reflected in the generated outputs. Addressing and mitigating these inherited biases remains a critical area for future research. The datasets used for evaluation are composed of publicly available videos from sources like Pexels, YouTube, and established academic benchmarks, respecting the data usage policies of these platforms.
Reproducibility statement
We are committed to ensuring the reproducibility of our research. All implementation details required to replicate our results are provided in the main paper and appendix. Our method is built upon the publicly available diffusion-pipe codebase, as cited in Section 4.1. We have released our specific source code, including scripts for mask-guided LoRA training and inference. The core of our experiments relies on publicly available pre-trained models, specifically Wan2.1-I2V and HunyuanVideo-I2V, which are clearly identified in Section 4.1. This section also details key hyperparameters, such as learning rates (1e-4), training steps (100+100), and hardware requirements (20GB GPU memory). Additional results and implementation strategies, including a low-memory approach, are described in the Appendix.
Appendix
Appendix A Implementation Details: Mask Acquisition GUI
To facilitate easy usage, we provide a Graphical User Interface (GUI) in our codebase (see Fig. 9). Users initialize the process by providing sparse clicks on the first frame to define the target object. The mask is then propagated automatically using SAM2.
Crucially, we utilize the bounding box derived from the segmentation for training, rather than the tight mask itself. As demonstrated in Section 4.3, this loose bounding box strategy is deliberate, providing a spatial buffer for natural transitions and structural changes. This workflow confirms that our method is highly automated and does not require manual frame-by-frame annotation.
Appendix B Additional Experimental Analysis
B.1 Input-Level vs. Feature-Level Masking
To verify the gain of our mask-aware fine-tuning strategy compared to feature-level strategies which utilize masks to constrain the background at the latent feature level, we conducted a comparison.
We established a feature-level baseline where the model is fine-tuned using the default Image-to-Video mask configuration (conditioning only on the first frame, prompting the model to learn global frame reconstruction). In this setup, background preservation is enforced as an inference-time intervention. Specifically, at each denoising timestep , we restrict the generation by mechanically blending the model’s predicted noisy latent with the noisy source latent using the binary segmentation mask (where 1 indicates the constrained background regions):
This strategy forces the background pixels to remain unchanged but ignores the semantic consistency between the edited region and the strictly constrained background.
The visual comparison in Figure 10 highlights the limitation of the feature-level constraint: foreground objects often appear detached. Quantitative results are reported in Table 3. Our input-conditional approach yields consistent gains across all key metrics, confirming that providing the spatiotemporal mask as a training input enables the model to synthesize more coherent and naturally integrated edits.
| Method | CLIP Score | DEQA Score | Input Similarity |
| Baseline (Feature-Level) | 0.8936 | 3.5878 | 0.7402 |
| Ours (Input-Level) | 0.9172 | 3.8013 | 0.7608 |
B.2 Effectiveness of Mask Scheduling for Appearance Learning
We further validate the role of the mask during the appearance learning stage. As demonstrated in Figure 11, using the standard I2V masking configuration (without our specific region-aware scheduling) fails to effectively update the object’s appearance.
This failure stems from the underlying mechanism of Image-to-Video models: the first frame is typically provided as a strong input condition. When fine-tuning on a reference image without masking, the model encounters a task where the input condition (the reference) is identical to the target output. Consequently, the optimization finds a trivial solution: it learns to simply copy the pixel information from the condition to the output. Because the appearance information is “leaked” through the input condition, the LoRA parameters do not internalize the visual attributes of the new object. Our proposed mask scheduling resolves this by masking the target region, thereby blocking this trivial shortcut and acting as a strict command that forces the LoRA to explicitly learn and generate the target appearance distribution.
B.3 Scalability Analysis
To verify the scalability and temporal stability of our method as requested, we evaluated the editing performance across video sequences of varying lengths: 5, 13, 21, and 49 frames. We utilize the CLIP Score as the metric, which measures the semantic alignment between generated frames and edited first frame by comparing their CLIP Radford et al. (2021) embedding similarity. As illustrated in Figure 12, our method demonstrates good stability.
Appendix C Efficiency Analysis & Low-Cost Strategy Details
C.1 Methodological Design: Temporal Windowing
The high VRAM requirement (20GB) of our standard training setting primarily stems from the necessity to process the full 49-frame sequence simultaneously during training. To address this, our “Low-Cost Strategy” introduces a training-time modification that circumvents this bottleneck without altering the inference pipeline. Specifically, we split the training video into shorter, overlapping sliding windows and update the LoRA weights based on local 9-frame segments rather than the entire sequence at once.
This design is theoretically grounded in the insight that motion LoRA primarily learns local dynamics. Since complex motions can be decomposed into continuous short-term patterns, training on 9-frame clips is sufficient to capture the necessary kinematics. Crucially, this fragmentation does not compromise global consistency because of our mask-aware conditioning. Since the unmasked background is fed as a visible input context, the model effectively learns “how the foreground moves relative to the fixed background” within each local window. This strong contextual anchoring ensures that the learned motion remains coherent and aligned with the environment when the full video is reassembled. As shown in Figure 13, the Low-Cost strategy maintains visual fidelity, motion smoothness, and temporal consistency comparable to standard full-frame training, validating the effectiveness of our decomposition approach.
C.2 Engineering Optimization: Swap Blocks
To further democratize access, we leveraged the “swap blocks” technique supported by the diffusion-pipe codebase222https://github.com/tdrussell/diffusion-pipe. This mechanism dynamically offloads frozen base model parameters to the CPU, keeping only trainable LoRA weights and active transformer blocks on the GPU. Table 4 quantifies the memory savings achieved by varying the blocks_to_swap parameter. By maximizing swapping (config 38), we reduce the peak VRAM to 8 GB, making the training pipeline accessible on consumer-grade GPUs.
| blocks_to_swap | 0 (None) | 16 | 32 (Default) | 38 (Max) |
| Peak VRAM (MB) | 25,097 | 16,305 | 9,987 | 7,617 |
C.3 Run-Time VRAM Comparison with Baselines
We benchmarked the run-time VRAM usage of our method against state-of-the-art inference-only baselines on a single NVIDIA RTX 4090 (, 49 frames).
| Method | AnyV2V | Go-with-the-flow* | Ours (Standard) | Ours (Low-Cost) |
| Peak VRAM | 13,680 MB | 4,398 / 21,512 MB | 21,522 MB | 7,617 MB |
As shown in Table 5, our optimized Low-Cost strategy achieves a memory footprint comparable to inference-only baselines, demonstrating broad hardware accessibility.
Appendix D Diverse Editing Results
We present visual results of our method on diverse editing tasks in Fig. 14 and Fig. 15, including object replacement, addition, and removal. We also include results for fast-motion and multi-person scenario in Fig. 16 and results for long-form video in Fig. 17.
Appendix E Results with HunyuanVideo-I2V Model
In addition to the main results based on Wan2.1-I2V, we also conducted experiments using HunyuanVideo-I2V (See Fig. 18). It demonstrates the generalization ability of our framework across different I2V architectures.
Appendix F Failure Cases
Figure 19 illustrates representative limitations encountered by our framework. We observe specific difficulties in text generation and preservation (as seen in the top rows). Even with mask constraints, the model often degrades high-frequency semantic symbols into blurred texture, likely due to the compression loss of the underlying VAE. Additionally, limitations arise in complex motion preservation, as demonstrated in the bottom rows where the character’s hair color is modified. In scenarios featuring rapid and highly stylized abstract motion, accurately decoupling the appearance update from the original dynamics becomes challenging, potentially resulting in unnatural deformations and disruptions to the structural integrity of the motion sequence.
Appendix G Details of User Study
To evaluate the performance of our method, we conducted a user study with 50 participants. Each participant was randomly shown 10 groups of results generated by different methods. For each group, the participants were asked to select the best result based on two criteria: reference similarity and visual quality. The user study interface is shown in Figure 20.
Instructions to Participants: Participants were provided with the following instructions:
”You will be shown several video sequences with different edits applied. Each sequence will contain three options, each generated using a different method. The methods will be presented in a random order to avoid bias. Your task is to choose the option that best matches the reference image in terms of visual quality and how well the edit aligns with the reference image. Pay close attention to the consistency of the edits across the frames and how naturally they interact with the environment.”
After viewing each video sequence, participants were asked to answer the following two questions for each video:
1.
Reference Similarity: Which video option best matches the reference image in terms of appearance and content?
2.
Visual Quality: Which video option has the highest overall visual quality, considering both the foreground edit and the background consistency?
The demographic information of the participants, including their age distribution and professional background, is shown in Figure 21. This figure demonstrates the diversity of the participant pool, ensuring a wide range of perspectives in the user study.
Appendix H Broader Impacts
The work presented in this paper introduces a flexible and controllable video editing method based on mask-aware LoRA fine-tuning. As with many advances in generative AI, our method holds both positive and negative societal implications.
Our video editing framework opens up new possibilities in creative fields, such as film production, digital art, and content creation, by enabling high-quality, real-time video edits with a high degree of control. This could streamline workflows for professionals in these industries, reducing the time and effort traditionally required for manual video editing. Additionally, our approach can be used in non-commercial domains, such as healthcare, where controllable video editing can assist in visualizing complex medical data, simulations, and surgical procedures. In such contexts, our method can aid in improving communication and understanding of visual information.
While the benefits are substantial, the potential for misuse of generative video models also exists. High-quality video editing tools can be misused to generate deepfakes, spreading misinformation or altering videos in malicious ways. Given the high realism of edited videos, there is a risk that such technologies could be exploited in political or social contexts, leading to challenges in verifying the authenticity of media. This raises concerns about privacy, security, and trust in video-based content.
To mitigate these risks, we advocate for responsible usage guidelines and the development of safety measures for content creation tools powered by generative models. For example, platforms using such technology could implement robust verification processes for user-generated content. Additionally, we encourage future research into model interpretability and the development of tools to detect manipulated media, ensuring that generated content is easily distinguishable from original footage.
Appendix I The Use of Large Language Models
During the preparation of this manuscript, we utilized a large language model (LLM) to aid in polishing the writing. Specifically, it was employed to improve grammar, correct spelling, enhance clarity, and ensure the conciseness of the text. It also assisted in refining sentence structures to better adhere to a formal academic style. All scientific contributions, including the core methodology, experimental design, results, and their interpretation, were solely conceived and articulated by the human authors. The authors have thoroughly reviewed and edited the final version of the text and take full responsibility for all content presented in this paper.